# On spectrum of corona product of duplication signed graph and its application

### v4

@@ 3.1 Structural properties
Structural property of corona of signed graphs has already being established by Adhikari at. el. [3], following that we present statistical information concerning the counts of vertices, edges, and traits in $\Gamma_{1}\circledast\Gamma_{2}$ for signed graphs $\Gamma_{1}=(G_{1},\sigma_{1},\mu_{1})$ and $\Gamma_{2}=(G_{2},\sigma_{2},\mu_{2})$. Clearly, the number of vertices in $\Gamma_{1}\circledast\Gamma_{2}$ is $|V|=|V_{1}|(2+|V_{2}|)$.
@@
The proofs of the formulas presented in Table 1and Table 2 directly follow from the definition of duplication add vertex corona product, and easily can be verified. Total triads of $\Gamma_{1}\circledast\Gamma_{2}$ is given by
@@
From Lemma 2.4, it is very much clear that $\Gamma_{1}\circledast\Gamma_{2}$ is unbalanced if $\Gamma_{2}$ is unbalanced. However, if $\Gamma_{2}$ is balanced then it is not certain that $\Gamma_{1}\circledast\Gamma_{2}$ is balanced. We can see the same in the Figure 2.
@@
The proof is the consequence of the observation that any positive or negative marked vertices of $\Gamma_{1}$, form a triad $T_{1}$ in $\Gamma_{1}\circledast\Gamma_{2}$ when there are edges of type $(i)$ and/or $(ii)$ in $\Gamma_{2}$. Otherwise, it results in the formation of a triad $T_{3}$ when $\Gamma_{2}$ includes an edge of type $(iii).$
@@ 3.2 Spectral properties
Cam McLeman, and Erin McNicholas [7] first introduced the coronal of graphs. Then Shu and Gui [8] generalised it and defined coronal for Laplacian and signless Laplacian matrix of unsigned graphs. Recently Amrik at. el. [15] defined coronal of signed graph as follows;
@@
[15] Given a signed graph $\Gamma=(G,\sigma,\mu)$ of order $n$ and $M$ be a graph matrix of $\Gamma$. Now, viewed as a matrix over the rational function field $\mathbb{C}(\lambda)$, the characteristic matrix $\lambda I_{n}-M$ has determinant non-zero, so it is invertible. The signed $M$-coronal $\chi_{M}(\lambda)\in\mathbb{C}(\lambda)$ of $\Gamma$ is defined as,
@@
$\big{(}\mu(\Gamma_{2}^{T})\otimes\phi(\Gamma_{1})\big{)}\big{(}(xI_{n_{2}}-M(\Gamma_{2}))^{-1}\otimes I_{n_{1}})\big{)}\big{(}\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})\big{)}=\chi_{M(\Gamma_{2})}(x)I_{n_{1}}\vspace{2pt}$ (2)
@@
where $p(A(\Gamma),\lambda)$ is a polynomial of degree $n-1$ and $f(M(\Gamma),\lambda)$ is the characteristics polynomial of the matrix $M$ of the signed graph $\Gamma$. If the greatest common divisor of these polynomials is a constant, then the above polynomial ratio can be further simplified as follows;
$=\frac{P_{d-1}(\lambda)}{F_{d}(\lambda)},$ (5)
where $P_{d-1}(\lambda)$ and $F_{d}(\lambda)$ are polynomials of degree $d-1$ and $d$ respectively and
@@
$W_{j}=\big{\{}v_{j}^{1},v_{j}^{2},\ldots,v_{j}^{n_{1}}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ j=1,2,\ldots,n_{2}$ (6)
Then $V(D\Gamma_{1})\cup W_{1}\cup W_{2}\cup\ldots\cup W_{n_{2}}$ is a partition of $V(\Gamma_{1}\circledast\Gamma_{2})$
@@
$d_{\Gamma_{1}\circledast\Gamma_{2}}(a_{i})=n_{2}+d_{\Gamma_{1}}(a_{i}),\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,n_{1}\$ $d_{\Gamma_{1}\circledast\Gamma_{2}}(u_{i})=d_{\Gamma_{1}}(u_{i}),\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,n_{1}\$ $d_{\Gamma_{1}\circledast\Gamma_{2}}(v^{i}_{j})=d_{\Gamma_{2}}(v_{j})+1,\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,n_{1},\leavevmode\nobreak\ \leavevmode\nobreak\ j=1,2,\ldots,n_{2}.$
@@ 3.2.1 A-spectra of signed graph
$\begin{split}S&=\begin{bmatrix}xI_{n_{1}}&-A(\Gamma_{1\mu})\\ -A(\Gamma_{1\mu})&xI_{n_{1}}\end{bmatrix}\begin{bmatrix}0\\ -\mu(\Gamma_{2})^{T}\otimes\phi(\Gamma_{1})\end{bmatrix}-\{(xI_{n_{2}}-A(\Gamma_{2}))^{-1}\otimes I_{n_{1}}\}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \begin{bmatrix}0&\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})\end{bmatrix}\end{split}$
@@
Also, by property of Kronecker product,
@@
If $\Gamma_{1}$ is balanced, then $\Gamma_{1}$ and $\leavevmode\nobreak\ \Gamma_{1\mu}$ are co-spectral and the characteristics polynomial of $\Gamma_{1}\circledast\Gamma_{2}$ is
@@
Let $\Gamma=(G,\sigma,\mu)$ be $k$-net-regular signed graph on $n$ nodes, then signed $A(\Gamma)-$coronal of $\Gamma$ is,
@@
Let $\Gamma_{1}$ be a signed graph with $n_{1}$ vertices and $\Gamma_{2}$ be a $k-$net regular signed graph with $n_{2}$ vertices. Also let the multiplicity of eigenvalue $k$ of $\Gamma_{2}$ be $p$, then the A-spectrum of $\Gamma_{1}\circledast\Gamma_{2}$ consist of,
  1. (i)
@@
Similarly, we use the following result by Bishal et.al. [17] in the following corollary.
@@
  1. (i)
    The eigenvalue 00 with multiplicity $n_{1}(n_{2}-1).$
  2. (ii)
@@
  1. (i)
    The eigenvalue 00 with multiplicity $n_{1}(n_{2}-1)$.
  2. (ii)
@@ 3.3 L-spectra of signed graph
Let $\Gamma_{1}$ be a $r_{1}-$regular signed graph with $n_{1}$ vertices and $\Gamma_{2}$ be any arbitrary signed graph on $n_{2}$ vertices with Laplacian spectrum $\gamma_{i1}\leq\gamma_{i2}\leq\ldots\leq\gamma_{in_{i}},\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2.$ Let $\Gamma_{1\mu}$ be the $\mu-$signed graph of $\Gamma_{1}$, and $\lambda^{\prime}_{11},\lambda^{\prime}_{12},\ldots,\lambda^{\prime}_{1n_{1}}$ be the eigenvalue of $\Gamma_{1\mu}.$ Then the Laplacian characteristics polynomial of duplication add vertex corona $\Gamma_{1}\circledast\Gamma_{2}$ is
@@
$\begin{split}f(L(\Gamma_{1}\circledast\Gamma_{2});x)&=\det\begin{bmatrix}(x-r_{1})I_{n_{1}}&A(\Gamma_{1\mu})&0_{n_{1}\times n_{1}n_{2}}\\ A(\Gamma_{1\mu})&(x-r_{1}-n_{2})I_{n_{1}}&\mu(\Gamma_{2})^{T}\otimes\phi(\Gamma_{1})\\ 0_{n_{1}n_{2}\times n_{1}}&\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})&((x-1)I_{n_{2}}-L(\Gamma_{2}))\otimes I_{n_{1}}\end{bmatrix}\\ &=\det\big{(}((x-1)I_{n_{2}}-L(\Gamma_{2}))\otimes I_{n_{1}}\big{)}\det(S),\\ \text{where }S&=\begin{bmatrix}(x-r_{1})I_{n_{1}}&A(\Gamma_{1\mu})\\ A(\Gamma_{1\mu})&(x-r_{1}-n_{2})I_{n_{1}}\end{bmatrix}-\begin{bmatrix}0\\ \mu(\Gamma_{2})^{T}\otimes\phi(\Gamma_{1})\end{bmatrix}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \big{(}((x-1)I_{n_{2}}-L(\Gamma_{2}))\otimes I_{n_{1}}\big{)}^{-1}\begin{bmatrix}0_{n_{1}n_{2}\times n_{1}}&\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})\end{bmatrix}\\ \end{split}$
@@
Also, by property of Kronecker product,
@@
Let $\Gamma_{1}$ be a $r_{1}-$regular signed graph with $n_{1}$ vertices and $\Gamma_{2}$ be any arbitrary signed graph on $n_{2}$ vertices with signless Laplacian spectrum $\nu_{i1}\leq\nu_{i2}\leq\ldots\leq\nu_{in_{i}},\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2.$ Let $\Gamma_{1\mu}$ be the $\mu-$signed graph of $\Gamma_{1}$, and $\lambda^{\prime}_{11},\lambda^{\prime}_{12},\ldots,\lambda^{\prime}_{1n_{1}}$ be the eigenvalue of $\Gamma_{1\mu}.$ Then the signless Laplacian characteristics polynomial of duplication add vertex corona $\Gamma_{1}\circledast\Gamma_{2}$ is
@@
The later part of the proof is similar to Theorem 3.4. ∎
@@ 4 Application
In this section we discuss some application of signed graph product $\Gamma_{1}\circledast\Gamma_{2}$ and its spectrum in generating family of integral signed graphs, and sequence of non-co-spectral equienergetic signed graphs.
@@ 4.1 Families of Integral Signed Graphs
A signed graph is called an integral signed graph if all the eigenvalues are integers. [18, 19]. The following Propositions shows the essential conditions for $\Gamma_{1}\circledast\Gamma_{2}$ to be an integral signed graph.
@@
Let $\Gamma_{1}$ be a signed graph with $n_{1}$ vertices and $\Gamma_{2}$ be a $k$-net-regular signed graph with $n_{2}$ vertices. Then $\Gamma_{1}\circledast\Gamma_{2}$ are integral signed graphs if and only if $\Gamma_{1\mu},\Gamma_{2}$ are integral signed graph and the roots of the equation, $x^{3}-kx^{2}-(n_{2}+{\lambda^{\prime}}_{1i}^{2})x+k{\lambda^{\prime}}_{1i}^{2}=0$, for each eigenvalue ${\lambda^{\prime}}_{1i}(i=1,2,\ldots,n_{1})$ of $\Gamma_{1\mu}$ are integers.
@@
Let $\Gamma_{1}$ be a signed graph with $n_{1}$ vertices and $\Gamma_{2}=(K_{1,n_{2}},\sigma_{2},\mu_{2})$ be a signed star on $n_{2}+1$ vertices with $V(\Gamma_{2})=\{v_{1},v_{2},\ldots,v_{n_{2}+1}\}$ where $d(v_{1})=n_{2}$. Then $\Gamma_{1}\circledast\Gamma_{2}$ are integral signed graphs if and only if $\Gamma_{1},\leavevmode\nobreak\ \Gamma_{2}$ are integral signed graph and the roots of the equation, $x^{4}-(2n_{2}+1+{\lambda}_{1i}^{2})x^{2}-2n_{2}\mu(v_{1})x+n_{2}{\lambda}_{1i}^{2}=0,$ for each eigenvalue ${\lambda}_{1i}(i=1,2,\ldots,n_{1})$ of $\Gamma_{1}$ are integers.

### v5

@@ 3.1 Structural properties
The structural property of the corona of signed graphs has already being established by Adhikari at el. [1], following that we present statistical information concerning the counts of vertices, edges, and traits in $\Gamma_{1}\circledast\Gamma_{2}$ for signed graphs $\Gamma_{1}=(G_{1},\sigma_{1},\mu_{1})$ and $\Gamma_{2}=(G_{2},\sigma_{2},\mu_{2})$. Clearly, the number of vertices in $\Gamma_{1}\circledast\Gamma_{2}$ is $|V|=|V_{1}|(2+|V_{2}|)$.
@@
The proofs of the formulas presented in Table 1and Table 2 directly follow from the definition of duplication add vertex corona product, and can easily be verified. Total triads of $\Gamma_{1}\circledast\Gamma_{2}$ is given by
@@
From Lemma 2.4, it is clear that $\Gamma_{1}\circledast\Gamma_{2}$ is unbalanced if $\Gamma_{2}$ is unbalanced. However, if $\Gamma_{2}$ is balanced, then it is not certain that $\Gamma_{1}\circledast\Gamma_{2}$ is balanced. We can see the same in the Figure 2.
@@
The proof is the consequence of the observation that any positive or negative marked vertices of $\Gamma_{1}$ form a triad $T_{1}$ in $\Gamma_{1}\circledast\Gamma_{2}$ when there are edges of type $(i)$ and/or $(ii)$ in $\Gamma_{2}$. Otherwise, it forms a triad $T_{3}$ when $\Gamma_{2}$ includes an edge of type $(iii).$
@@ 3.2 Spectral properties
Cam McLeman and Erin McNicholas [11] first introduced the coronal of graphs. Then Shu and Gui [5] generalized it and defined coronal for Laplacian and signless Laplacian matrix of unsigned graphs. Recently, Amrik at. el. [17] defined the coronal of the signed graph as follows;
@@
[17] Given a signed graph $\Gamma=(G,\sigma,\mu)$ of order $n$ and $M$ be a graph matrix of $\Gamma$. Now, viewed as a matrix over the rational function field $\mathbb{C}(\lambda)$, the characteristic matrix $\lambda I_{n}-M$ has a non-zero determinant, which is invertible. The signed $M$-coronal $\chi_{M}(\lambda)\in\mathbb{C}(\lambda)$ of $\Gamma$ is defined as,
@@
$\big(\mu(\Gamma_{2}^{T})\otimes\phi(\Gamma_{1})\big)\big((xI_{n_{2}}-M(\Gamma_{2}))^{-1}\otimes I_{n_{1}})\big)\big(\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})\big)=\chi_{M(\Gamma_{2})}(x)I_{n_{1}}\vskip 2.0pt$ (2)
@@
where $p(A(\Gamma),\lambda)$ is a polynomial of degree n1n-1 and $f(M(\Gamma),\lambda)$ is the characteristics polynomial of the matrix $M$ of the signed graph $\Gamma$. If the greatest common divisor of these polynomials is a constant, then the above polynomial ratio can be further simplified as follows:
$\frac{P_{d-1}(\lambda)}{F_{d}(\lambda)},$ (5)
where $P_{d-1}(\lambda)$ and $F_{d}(\lambda)$ are polynomials of degree d1d-1 and $d$ respectively and
@@
$W_{j}=\big\{v_{j}^{1},v_{j}^{2},\ldots,v_{j}^{n_{1}}\big\},\penalty 10000\ \penalty 10000\ \penalty 10000\ j=1,2,\ldots,n_{2}$ (6)
Then $V(D\Gamma_{1})\cup W_{1}\cup W_{2}\cup\ldots\cup W_{n_{2}}$ is a partition of $V(\Gamma_{1}\circledast\Gamma_{2})$ .
@@
$d_{\Gamma_{1}\circledast\Gamma_{2}}(a_{i})=n_{2}+d_{\Gamma_{1}}(a_{i}),\penalty 10000\ \penalty 10000\ i=1,2,\ldots,n_{1}\\$ $d_{\Gamma_{1}\circledast\Gamma_{2}}(u_{i})=d_{\Gamma_{1}}(u_{i}),\penalty 10000\ \penalty 10000\ i=1,2,\ldots,n_{1}\\$ $d_{\Gamma_{1}\circledast\Gamma_{2}}(v^{i}_{j})=d_{\Gamma_{2}}(v_{j})+1,\penalty 10000\ \penalty 10000\ i=1,2,\ldots,n_{1},\penalty 10000\ \penalty 10000\ j=1,2,\ldots,n_{2}.$
@@ 3.2.1 A-spectra of signed graph
$\begin{split}S&=\begin{bmatrix}xI_{n_{1}}&-A(\Gamma_{1\mu})\\ -A(\Gamma_{1\mu})&xI_{n_{1}}\end{bmatrix}\begin{bmatrix}0\\ -\mu(\Gamma_{2})^{T}\otimes\phi(\Gamma_{1})\end{bmatrix}-\{(xI_{n_{2}}-A(\Gamma_{2}))^{-1}\otimes I_{n_{1}}\}\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \begin{bmatrix}0&\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})\end{bmatrix}\end{split}$
@@
Also, by the property of the Kronecker product,
@@
If $\Gamma_{1}$ is balanced, then $\Gamma_{1}$ and $\penalty 10000\ \Gamma_{1\mu}$ are co-spectral and the characteristics polynomial of $\Gamma_{1}\circledast\Gamma_{2}$ is
@@
Let $\Gamma=(G,\sigma,\mu)$ be a non-null, $(r,k)$-co-regular signed graph on $n$ nodes, then signed $A(\Gamma)$-coronal of $\Gamma$ is,
@@
Let $\Gamma_{1}$ be a signed graph with $n_{1}$ vertices, and $\Gamma_{2}$ be a $(r,k)$-co-regular signed graph with $n_{2}$ vertices. Also let the multiplicity of eigenvalue $k$ of $\Gamma_{2}$ be $p$, then the A-spectrum of $\Gamma_{1}\circledast\Gamma_{2}$ consist of,
  1. 1.
@@
Similarly, we use the following result by Bishal et al. [18] in the following corollary.
@@
  1. 1.
    The eigenvalue 0 with multiplicity $n_{1}(n_{2}-1).$
  2. 2.
@@
  1. 1.
    The eigenvalue 0 with multiplicity $n_{1}(n_{2}-1)$.
  2. 2.
@@ 3.3 L-spectra of signed graph
Let $\Gamma_{1}$ be a $r_{1}$-regular signed graph with $n_{1}$ vertices and $\Gamma_{2}$ be any arbitrary signed graph on $n_{2}$ vertices with Laplacian spectrum $\gamma_{i1}\leq\gamma_{i2}\leq\ldots\leq\gamma_{in_{i}},\penalty 10000\ \penalty 10000\ i=1,2.$ Let $\Gamma_{1\mu}$ be the $\mu$-signed graph of $\Gamma_{1}$, and $\lambda^{\prime}_{11},\lambda^{\prime}_{12},\ldots,\lambda^{\prime}_{1n_{1}}$ be the eigenvalue of $\Gamma_{1\mu}.$ Then the Laplacian characteristics polynomial of duplication add vertex corona $\Gamma_{1}\circledast\Gamma_{2}$ is
@@
$\begin{split}f(L(\Gamma_{1}\circledast\Gamma_{2});x)&=\det\begin{bmatrix}(x-r_{1})I_{n_{1}}&A(\Gamma_{1\mu})&0_{n_{1}\times n_{1}n_{2}}\\ A(\Gamma_{1\mu})&(x-r_{1}-n_{2})I_{n_{1}}&\mu(\Gamma_{2})^{T}\otimes\phi(\Gamma_{1})\\ 0_{n_{1}n_{2}\times n_{1}}&\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})&((x-1)I_{n_{2}}-L(\Gamma_{2}))\otimes I_{n_{1}}\end{bmatrix}\\ &=\det\big(((x-1)I_{n_{2}}-L(\Gamma_{2}))\otimes I_{n_{1}}\big)\det(S),\\ \text{where }S&=\begin{bmatrix}(x-r_{1})I_{n_{1}}&A(\Gamma_{1\mu})\\ A(\Gamma_{1\mu})&(x-r_{1}-n_{2})I_{n_{1}}\end{bmatrix}-\begin{bmatrix}0\\ \mu(\Gamma_{2})^{T}\otimes\phi(\Gamma_{1})\end{bmatrix}\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \big(((x-1)I_{n_{2}}-L(\Gamma_{2}))\otimes I_{n_{1}}\big)^{-1}\begin{bmatrix}0_{n_{1}n_{2}\times n_{1}}&\mu(\Gamma_{2})\otimes\phi(\Gamma_{1})\end{bmatrix}\\ \end{split}$
@@
Also, by the property of the Kronecker product,
@@
Let $\Gamma_{1}$ be a $r_{1}$-regular signed graph with $n_{1}$ vertices and $\Gamma_{2}$ be any arbitrary signed graph on $n_{2}$ vertices with signless Laplacian spectrum $\nu_{i1}\leq\nu_{i2}\leq\ldots\leq\nu_{in_{i}},\penalty 10000\ \penalty 10000\ i=1,2.$ Let $\Gamma_{1\mu}$ be the $\mu$-signed graph of $\Gamma_{1}$, and $\lambda^{\prime}_{11},\lambda^{\prime}_{12},\ldots,\lambda^{\prime}_{1n_{1}}$ be the eigenvalue of $\Gamma_{1\mu}.$ Then the signless Laplacian characteristics polynomial of duplication add vertex corona $\Gamma_{1}\circledast\Gamma_{2}$ is
@@
The latter part of the proof is similar to Theorem 3.4. ∎
@@ 4 Application
This section discusses some applications of the signed graph product $\Gamma_{1}\circledast\Gamma_{2}$ and its spectrum in generating a family of integral signed graphs and a sequence of non-co-spectral equienergetic signed graphs.
@@ 4.1 Families of Integral Signed Graphs
A signed graph is called an integral signed graph if all the eigenvalues are integers. [2, 10]. The following Propositions show the essential conditions for $\Gamma_{1}\circledast\Gamma_{2}$ to be an integral signed graph.
@@
Let $\Gamma_{1}$ be a signed graph with $n_{1}$ vertices, and $\Gamma_{2}$ be a $k$-net-regular signed graph with $n_{2}$ vertices. Then $\Gamma_{1}\circledast\Gamma_{2}$ are integral signed graphs if and only if $\Gamma_{1\mu},\Gamma_{2}$ are integral signed graph and the roots of the equation, $x^{3}-kx^{2}-(n_{2}+{\lambda^{\prime}}_{1i}^{2})x+k{\lambda^{\prime}}_{1i}^{2}=0$, for each eigenvalue ${\lambda^{\prime}}_{1i}(i=1,2,\ldots,n_{1})$ of $\Gamma_{1\mu}$ are integers.
@@
Let $\Gamma_{1}$ be a signed graph with $n_{1}$ vertices and $\Gamma_{2}=(K_{1,n_{2}},\sigma_{2},\mu_{2})$ be a signed star on $n_{2}+1$ vertices with $V(\Gamma_{2})=\{v_{1},v_{2},\ldots,v_{n_{2}+1}\}$ where $d(v_{1})=n_{2}$. Then $\Gamma_{1}\circledast\Gamma_{2}$ are integral signed graphs if and only if $\Gamma_{1},\penalty 10000\ \Gamma_{2}$ are integral signed graph and the roots of the equation, $x^{4}-(2n_{2}+1+{\lambda}_{1i}^{2})x^{2}-2n_{2}\mu(v_{1})x+n_{2}{\lambda}_{1i}^{2}=0,$ for each eigenvalue ${\lambda}_{1i}(i=1,2,\ldots,n_{1})$ of $\Gamma_{1}$ are integers.